write the equation in the form Ax+By+C=0
a) it is a vertical line which passes through (-3,2)
b) it is a horizontal line passing through (4,-1)
c) parallel line parallel to 3x+2y=4 and passes through (4,-1)

Answer by lwsshak3(11628)

write the equation in the form Ax+By+C=0
..
a) it is a vertical line which passes through (-3,2)
equation: x+3=0
..
b) it is a horizontal line passing through (4,-1)
equation: y+1=0
..
c) parallel line parallel to 3x+2y=4 and passes through (4,-1)
Standard form of equation for a straight line: y=mx+b, m=slope, b=y-intercept
3x+2y=4
2y=-3x+4
y=-3x/2+2
slope,m=-3/2
equation of parallel line:
y=-3x/2+b
solving for b using coordinates of given point(4,-1) which is on the line.
-1=-3*4/2+b
-1=-6+b
b=5
equation of parallel line: y=-3x/2+5